Sharkattack123:

The function f(x) = x + 5 represents the length of a rectangle, and the function g(x) = 7x - 2 represents the width of the rectangle. What is the area of the rectangle if x = 3 A. 56 B. 27 C. 285 D. 152

Laylalyssa:

Okay so it tells us that x=3 so we have to plug that in for f(x) and g(x) f(x)=3+5 f(x)=8 g(x)=7(3)-2 g(x)=19 Area of a rectangle is Length times width 8 times 19 is 152 so 152 is the answer

Sharkattack123:

Thank you 😭

Laylalyssa:

yww<3